Find all numbers such that the square of the number is 48 more than double the number. If you find more than one, then list all the numbers you find in increasing order, separated by commas.
Answer by ankor@dixie-net.com(22740)

Find all numbers such that the square of the number is 48 more than double the number.
If you find more than one, then list all the numbers you find in increasing order, separated by commas.
:
n^2 = 2n + 48
n^2 - 2n - 48 = 0; a quadratic equation
Factor to
(n-8)(n+6) = 0
two solutions
n = -6, 8
:
you can check both solutions in the original statement